# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ 6.4 A Dual-Constraint Framework: RCC and the Exact Quantum Speed Limit
The core contribution of this work is the establishment of a rigorous information-theoretic lower bound (Theorem 3.1) for any quantum process that prepares a given state, derived from the structured information of the final state. This bound is essentially a static constraint, uniquely determined by the information content of the final state $\rho$ relative to its structured vacuum $\sigma_{R}$, and is independent of the specific evolution path. A natural question follows: does a complementary constraint exist, one that originates from the process dynamics itself, to limit the efficiency of the evolution? Quantum Speed Limits (QSLs) provide the theoretical framework for this. Recent advancements, particularly the "exact-QSL," have established an upper bound on this speed, determined by the variance of the non-classical part of the generator relative to the instantaneous state—a bound that is tighter than the Mandelstam–Tamm bound [43] [44]. In this section, we juxtapose the RCC information-theoretic lower bound with the exact-QSL dynamical upper bound to construct a dual-constraint theoretical framework for quantum processes, revealing the complementary relationship between information cost and dynamical efficiency. We first consider a pure-state unitary evolution process driven by a Hamiltonian $H$, evolving from an initial state $|\psi_{0}\rangle$ to a final state $\rho_{T}:=|\psi_{T}\rangle\langle\psi_{T}|$. Following Hall’s exact uncertainty relation [45], the generator can be decomposed relative to the instantaneous state $\rho_{t}$ as $H=H_{\text{cl}}+H_{\text{nc}}$, where only the uncertainty of the non-classical part, $\Delta H_{\text{nc}}$, contributes to the instantaneous speed of the quantum state ray [44]. The energy uncertainty of an operator $X$ is defined as $\Delta X(t)=\sqrt{\langle X^{2}\rangle_{t}-\langle X\rangle_{t}^{2}}$, and the time average is defined as $\langle f\rangle_{T}:=T^{-1}\int_{0}^{T}[f(t)]\,dt$.
@@
Definition 6.4.1 (Fubini-Study Angle). For an initial state $|\psi_{0}\rangle$ and a final state $|\psi_{T}\rangle$, the Fubini-Study angle $\theta$ between them is defined as[44]:
$\theta=\arccos(|\langle\psi_{0}|\psi_{T}\rangle|)\in[0,\pi/2]$ (6.4)
@@
Theorem 6.4.1 (exact-QSL, adapted from [44]). For any pure-state unitary evolution, its Fubini-Study angle $\theta$ is constrained by the following dynamical upper bound:
$\theta\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H(t)]\,dt$ (6.5)
@@
$T=\hbar\frac{l(\phi_{t})|_{0}^{T}}{\langle\Delta H_{\text{nc}}\rangle_{T}}$ (6.6)
where $l(\phi_{t})$ is the Wootters/statistical distance [46], and $\theta\leq l(\phi_{t})|_{0}^{T}$. In cases such as an effective two-level system where the survival probability $p_{t}=|\langle\psi_{0}|\psi_{t}\rangle|^{2}$ is monotonic, or when the Hamiltonian $H^{2}=I$ (self-inverse), the equality $\theta=(1/\hbar)\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt$ holds. [^4] For a single-qubit system with a self-inverse Hamiltonian $H=\sigma_{x}$ (taking $\hbar=1$) and initial state $|\psi_{0}\rangle=|0\rangle$, the evolved state is $|\psi_{t}\rangle=\cos(t)|0\rangle-i\sin(t)|1\rangle$. Here, $\theta(t)=t$ and $\Delta H_{\text{nc}}(t)=1$, so the equality in (6.5) holds, providing a minimal, achievable example of this theory. We can now juxtapose this dynamical ceiling with the RCC information-theoretic floor from this work. The two bounds apply to different objects (circuit cost vs. state geometric distance) and do not form a sandwich bound on the same quantity.
@@
$C_{\text{opt}}^{(\epsilon)}(\rho_{T})\geq\frac{D(\rho_{T}\|\sigma_{R})}{\log(\Gamma_{R})}+\dots$ (6.7)
@@
$\theta\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt$ (6.8)
@@
$\text{RCC}_{R}(\rho_{T})\lesssim C_{\text{FS}}(\rho_{T})=\theta$ (6.10)
The $\lesssim$ symbol in Eq.(6.10) emphasizes that this is a heuristic alignment, not a universal proposition. This relation suggests that a state with higher information-structural complexity (large RCC) may also need to traverse a greater geometric distance to be reached from a simple initial state. It is crucial to contextualize this heuristic. In other frameworks, such as Nielsen’s geometric complexity which employs a Finsler metric with penalty factors, the relationship is often reversed, yielding a model-dependent lower bound of the form $C_{\text{Nielsen}}\geq\lambda_{\gamma}\theta$ [47]. Our heuristic alignment, therefore, specifically pertains to the comparison between the information content (RCC) and the minimal geometric path ($\theta$). With this understanding, we adopt a conservative formulation for experimental purposes, proposing the hypothesis $\theta\geq c_{h}\,RCC_{R}(\rho_{T})$, where $0<c_{h}\leq 1$ is a model-dependent heuristic constant. Combining this conservative hypothesis with the traditional QSL bound $T\geq\hbar\theta/\langle\Delta H\rangle_{T}$ leads to a meaningful physical conclusion.
@@
$T\geq\frac{\hbar}{\langle\Delta H\rangle_{T}}\cdot\theta\geq\frac{\hbar c_{h}}{\langle\Delta H\rangle_{T}}\cdot\left[\frac{D(\rho_{T}\|\sigma_{R})}{\log(\Gamma_{R})}+\dots\right]$ (6.11)
@@
This result connects the final state’s RCC information threshold with the process’s time-averaged energy uncertainty $\langle\Delta H\rangle_{T}$, collectively providing an auditable lower bound on the minimum evolution time. The dual-constraint framework proposed in this section unifies the static constraints of information theory and the dynamic constraints of dynamics, offering a more comprehensive and profound perspective for evaluating the resource consumption and evolutionary efficiency of quantum processes. It suggests that a deeper intrinsic unity may exist between computation, information, and dynamics than previously thought. Future open questions include: on which families of physically realizable systems does a stable quantitative relationship exist between RCC and $\theta$, and how to consistently generalize this framework to open quantum systems [48, 49].
6.5 Outlook for Future Work
@@
The establishment of the RCC theoretical framework opens up avenues for future exploration in several directions. Firstly, the generalization from state complexity to the complexity of unitary operators/channels is a natural direction. By applying the idea of reference sets to Liouville space, we might be able to define an RCC for a unitary operator or a quantum channel, to quantify the intrinsic complexity of realizing a specific quantum operation compared to "free" operations (such as operations that preserve a specific subspace). Secondly, systematically connecting RCC with open quantum systems and quantum thermodynamics is another promising direction. For instance, by modeling processes that preserve energy eigen-populations or pinching processes of specific configurations as R-preserving channels, one can use the monotonicity of RCC, Eq. (5.1), to provide a non-trivial upper bound on the rate of information dissipation. Comparing this bound quantitatively with the change in thermodynamic quantities such as entropy production rate or non-equilibrium free energy is expected to lead to a series of testable inequalities connecting information-theoretic costs with thermodynamic prices. Finally, deepening the application of RCC in holographic duality will be an exciting long-term goal. As described in Sec. 5.3, RCC can serve as a structured calibration benchmark to test the order-preserving capability and proportional stability of different geometric candidate quantities (such as the CV/CA conjectures) across multiple families of states. If inconsistencies are found, they may suggest the existence of structural biases in the duality proposal and provide clues for finding directions for correction. (We provide a rigorous framework in Appendix F.2 for calculating a controllable, conservative lower bound on complexity in the presence of energy or mode cutoffs.)
7 Conclusion
@@
This paper has systematically proposed and constructed the theoretical framework of “Reference-Contingent Complexity” ($RCC$), aiming to provide a self-consistent and operational solution for addressing the long-standing “structural fairness problem” in quantum information science. Starting from first principles, we formalize the intrinsic structure of a system—whether it originates from symmetries, conservation laws, or deliberate encoding—as an a priori “reference set,” and, using this as a benchmark, we define complexity as the information gap of a quantum state relative to this structured background, a gap which can be measured by relative entropy. This construction fundamentally ensures that our measure only prices the non-trivial quantum information that must be generated at the cost of computational resources. The core theoretical contribution of this paper is the proof of a main theorem (Theorem 3.1), which establishes that $RCC$ is a rigorous, theorem-level lower bound on the optimal number of atomic instruction steps for any universal quantum circuit. This theorem is not only mathematically rigorous, but its physical meaning also reveals a profound connection between the information-theoretic properties of a state (the information gap $D(\rho||\sigma_{R})$ and the spectral skew $\Delta_{\mathrm{spec}}$) and the computational cost of the dynamical process. To ensure the physical relevance of this theory, we have further proposed a set of operational principles based on quantum hypothesis testing, projective witnesses, and dephasing entropy bounds (detailed in Chapter 4), demonstrating that the lower bound from $RCC$ can be estimated through experiments and numerical simulations, thereby connecting an abstract theoretical concept with concretely measurable physical observables. As a theoretical tool, the value of $RCC$ is manifested in its broad application interfaces. We have shown the well-behaved mathematical properties it possesses as a proper resource measure, such as monotonicity under free operations (Sec. 5.1). We have elucidated how it complements and deepens the theoretical framework for the performance frontier of quantum dynamics, elevating a process-dependent complexity measure to a state function determined solely by the final state, as shown in (Sec. 5.2). Furthermore, we propose that the structural fairness of $RCC$ makes it a promising candidate to play a key role as a calibration benchmark in the exploration of fundamental physical questions such as the “complexity-geometry” holographic duality (Sec. 5.3). In summary, Reference-Contingent Complexity provides a foundational and structurally fair theoretical tool for understanding and quantifying the physical costs of quantum computation. It offers a unified “meter stick” for us to compare and diagnose increasingly complex quantum algorithms and hardware platforms, and opens up a solid and operational new path for exploring the deep unification of information, dynamics, and spacetime geometry. We are well aware that the establishment of any theoretical framework is merely the beginning of an exploration, not its end. Many theoretical properties of $RCC$, such as the tightness of its lower bound (Conjecture 6.1), its generalization to systems with non-Abelian symmetries, and its quantitative connection to quantum chaos and thermodynamics, remain open questions full of opportunities. We hope that this work will inspire more theoretical and experimental explorations within the field, jointly advancing towards a deeper understanding of the fundamental concept of “complexity” in the quantum world.
@@ A.1 Formalization of the Computational Model
(Note: Formalization of the computational model. For the purpose of the rigorous mathematical proofs that follow, this appendix adopts the computational model already established in the main text. The physical meaning and conventions for the optimal circuit complexity $C_{opt}$ and the information bandwidth parameter $\Gamma_{R}$ (particularly concerning how to handle parallel circuits and continuous-parameter gates) have been fully and unambiguously specified in the notes of Section 2.1 of the main text.)
@@ A.2 Encoding Upper Bound: From Optimal Instruction Steps to Kolmogorov Complexity
$K_{\varepsilon}(\rho|R)\leq C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\cdot\log\Gamma_{R}+c_{a}\cdot\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+c_{b}\cdot\log(1/\varepsilon)+O(1)$ (A.2)
@@
(Prefix Code Specification. A self-describing prefix code (such as Elias-delta) is used for $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ and $\varepsilon$. Therefore, there exist constants $c_{a},c_{b}=O(1)$ that depend only on the chosen universal machine and the prefix code, and are independent of $\rho,\Gamma_{R},d_{R}$, such that Eq. (A.2) holds.)

### v3

@@ VI.4 A Dual-Constraint Framework: RCC and the Exact Quantum Speed Limit
The core contribution of this work is the establishment of a rigorous information-theoretic lower bound (Theorem 3.1) for any quantum process that prepares a given state, derived from the structured information of the final state. This bound is essentially a static constraint, uniquely determined by the information content of the final state $\rho$ relative to its structured vacuum $\sigma_{R}$, and is independent of the specific evolution path. A natural question follows: does a complementary constraint exist, one that originates from the process dynamics itself, to limit the efficiency of the evolution? Quantum Speed Limits (QSLs) provide the theoretical framework for this. Recent advancements, particularly the "exact-QSL," have established an upper bound on this speed, determined by the variance of the non-classical part of the generator relative to the instantaneous state—a bound that is tighter than the Mandelstam–Tamm bound [MandelstamTamm1945] [pati2023exact]. In this section, we juxtapose the RCC information-theoretic lower bound with the exact-QSL dynamical upper bound to construct a dual-constraint theoretical framework for quantum processes, revealing the complementary relationship between information cost and dynamical efficiency. We first consider a pure-state unitary evolution process driven by a Hamiltonian $H$, evolving from an initial state $|\psi_{0}\rangle$ to a final state $\rho_{T}:=|\psi_{T}\rangle\langle\psi_{T}|$. Following Hall’s exact uncertainty relation [Hall2001], the generator can be decomposed relative to the instantaneous state $\rho_{t}$ as $H=H_{\text{cl}}+H_{\text{nc}}$, where only the uncertainty of the non-classical part, $\Delta H_{\text{nc}}$, contributes to the instantaneous speed of the quantum state ray [pati2023exact]. The energy uncertainty of an operator $X$ is defined as $\Delta X(t)=\sqrt{\langle X^{2}\rangle_{t}-\langle X\rangle_{t}^{2}}$, and the time average is defined as $\langle f\rangle_{T}:=T^{-1}\int_{0}^{T}[f(t)]\,dt$.
@@
Definition 6.4.1 (Fubini-Study Angle). For an initial state $|\psi_{0}\rangle$ and a final state $|\psi_{T}\rangle$, the Fubini-Study angle $\theta$ between them is defined as:
$\theta=\arccos(|\langle\psi_{0}|\psi_{T}\rangle|)\in[0,\pi/2]$ (23)
@@
Theorem 6.4.1 (exact-QSL, adapted from [pati2023exact]). For any pure-state unitary evolution, its Fubini-Study angle $\theta$ is constrained by the following dynamical upper bound:
$\theta\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H(t)]\,dt$ (24)
@@
$T=\hbar\frac{l(\phi_{t})|_{0}^{T}}{\langle\Delta H_{\text{nc}}\rangle_{T}}$ (25)
where $l(\phi_{t})$ is the Wootters/statistical distance [Wootters81], and $\theta\leq l(\phi_{t})|_{0}^{T}$. In cases such as an effective two-level system where the survival probability $p_{t}=|\langle\psi_{0}|\psi_{t}\rangle|^{2}$ is monotonic, or when the Hamiltonian $H^{2}=I$ (self-inverse), the equality $\theta=(1/\hbar)\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt$ holds. [^4] For a single-qubit system with a self-inverse Hamiltonian $H=\sigma_{x}$ (taking $\hbar=1$) and initial state $|\psi_{0}\rangle=|0\rangle$, the evolved state is $|\psi_{t}\rangle=\cos(t)|0\rangle-i\sin(t)|1\rangle$. Here, $\theta(t)=t$ and $\Delta H_{\text{nc}}(t)=1$, so the equality in (24) holds, providing a minimal, achievable example of this theory. We can now juxtapose this dynamical ceiling with the RCC information-theoretic floor from this work. The two bounds apply to different objects (circuit cost vs. state geometric distance) and do not form a sandwich bound on the same quantity.
@@
$C_{\text{opt}}^{(\epsilon)}(\rho_{T})\geq\frac{D(\rho_{T}\|\sigma_{R})}{\log(\Gamma_{R})}+\dots$ (26)
@@
$\theta\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt$ (27)
@@
$\text{RCC}_{R}(\rho_{T})\lesssim C_{\text{FS}}(\rho_{T})=\theta$ (29)
The $\lesssim$ symbol in Eq.(29) emphasizes that this is a heuristic alignment, not a universal proposition. This relation suggests that a state with higher information-structural complexity (large RCC) may also need to traverse a greater geometric distance to be reached from a simple initial state. It is crucial to contextualize this heuristic. In other frameworks, such as Nielsen’s geometric complexity which employs a Finsler metric with penalty factors, the relationship is often reversed, yielding a model-dependent lower bound of the form $C_{\text{Nielsen}}\geq\lambda_{\gamma}\theta$ [NielsenGeometry06]. Our heuristic alignment, therefore, specifically pertains to the comparison between the information content (RCC) and the minimal geometric path ($\theta$). With this understanding, we adopt a conservative formulation for experimental purposes, proposing the hypothesis $\theta\geq c_{h}\,RCC_{R}(\rho_{T})$, where $0<c_{h}\leq 1$ is a model-dependent heuristic constant. Combining this conservative hypothesis with the traditional QSL bound $T\geq\hbar\theta/\langle\Delta H\rangle_{T}$ leads to a meaningful physical conclusion.
@@
$T\geq\frac{\hbar}{\langle\Delta H\rangle_{T}}\cdot\theta\geq\frac{\hbar c_{h}}{\langle\Delta H\rangle_{T}}\cdot\left[\frac{D(\rho_{T}\|\sigma_{R})}{\log(\Gamma_{R})}+\dots\right]$ (30)
@@
This result connects the final state’s RCC information threshold with the process’s time-averaged energy uncertainty $\langle\Delta H\rangle_{T}$, collectively providing an auditable lower bound on the minimum evolution time. The dual-constraint framework proposed in this section unifies the static constraints of information theory and the dynamic constraints of dynamics, offering a more comprehensive and profound perspective for evaluating the resource consumption and evolutionary efficiency of quantum processes. It suggests that a deeper intrinsic unity may exist between computation, information, and dynamics than previously thought. Future open questions include: on which families of physically realizable systems does a stable quantitative relationship exist between RCC and $\theta$, and how to consistently generalize this framework to open quantum systems [DeffnerCampbell17, TaddeiPRL13].
VI.5 Outlook for Future Work
@@
The establishment of the RCC theoretical framework opens up avenues for future exploration in several directions. Firstly, the generalization from state complexity to the complexity of unitary operators/channels is a natural direction. By applying the idea of reference sets to Liouville space, we might be able to define an RCC for a unitary operator or a quantum channel, to quantify the intrinsic complexity of realizing a specific quantum operation compared to "free" operations (such as operations that preserve a specific subspace). Secondly, systematically connecting RCC with open quantum systems and quantum thermodynamics is another promising direction. For instance, by modeling processes that preserve energy eigen-populations or pinching processes of specific configurations as R-preserving channels, one can use the monotonicity of RCC, Eq. (12), to provide a non-trivial upper bound on the rate of information dissipation. Comparing this bound quantitatively with the change in thermodynamic quantities such as entropy production rate or non-equilibrium free energy is expected to lead to a series of testable inequalities connecting information-theoretic costs with thermodynamic prices. Finally, deepening the application of RCC in holographic duality will be an exciting long-term goal. As described in Sec. V.3, RCC can serve as a structured calibration benchmark to test the order-preserving capability and proportional stability of different geometric candidate quantities (such as the CV/CA conjectures) across multiple families of states. If inconsistencies are found, they may suggest the existence of structural biases in the duality proposal and provide clues for finding directions for correction. (We provide a rigorous framework in Appendix F.2 for calculating a controllable, conservative lower bound on complexity in the presence of energy or mode cutoffs.)
VII Conclusion
@@
This paper has systematically proposed and constructed the theoretical framework of “Reference-Contingent Complexity” ($RCC$), aiming to provide a self-consistent and operational solution for addressing the long-standing “structural fairness problem” in quantum information science. Starting from first principles, we formalize the intrinsic structure of a system—whether it originates from symmetries, conservation laws, or deliberate encoding—as an a priori “reference set,” and, using this as a benchmark, we define complexity as the information gap of a quantum state relative to this structured background, a gap which can be measured by relative entropy. This construction fundamentally ensures that our measure only prices the non-trivial quantum information that must be generated at the cost of computational resources. The core theoretical contribution of this paper is the proof of a main theorem (Theorem 3.1), which establishes that $RCC$ is a rigorous, theorem-level lower bound on the optimal number of atomic instruction steps for any universal quantum circuit. This theorem is not only mathematically rigorous, but its physical meaning also reveals a profound connection between the information-theoretic properties of a state (the information gap $D(\rho||\sigma_{R})$ and the spectral skew $\Delta_{\mathrm{spec}}$) and the computational cost of the dynamical process. To ensure the physical relevance of this theory, we have further proposed a set of operational principles based on quantum hypothesis testing, projective witnesses, and dephasing entropy bounds (detailed Sec. IV), demonstrating that the lower bound from $RCC$ can be estimated through experiments and numerical simulations, thereby connecting an abstract theoretical concept with concretely measurable physical observables. As a theoretical tool, the value of $RCC$ is manifested in its broad application interfaces. We have shown the well-behaved mathematical properties it possesses as a proper resource measure, such as monotonicity under free operations (Sec. V.1). We have elucidated how it complements and deepens the theoretical framework for the performance frontier of quantum dynamics, elevating a process-dependent complexity measure to a state function determined solely by the final state, as shown in (Sec. V.2). Furthermore, we propose that the structural fairness of $RCC$ makes it a promising candidate to play a key role as a calibration benchmark in the exploration of fundamental physical questions such as the “complexity-geometry” holographic duality (Sec. V.3). In summary, Reference-Contingent Complexity provides a structurally fair theoretical toolfor understanding and quantifying the physical costs of quantum computation. It offers auseful benchmark for comparing and diagnosing complex quantum algorithms and hardware platforms, and suggests a concrete route for exploring possible links between information, dynamics, and spacetime geometry.
@@ A.1 Formalization of the Computational Model
(Note: Formalization of the computational model. For the purpose of the rigorous mathematical proofs that follow, this appendix adopts the computational model already established in the main text. The physical meaning and conventions for the optimal circuit complexity $C_{opt}$ and the information bandwidth parameter $\Gamma_{R}$ (particularly concerning how to handle parallel circuits and continuous-parameter gates) have been fully and unambiguously specified in the notes of Section II.1 of the main text.)
@@ A.2 Encoding Upper Bound: From Optimal Instruction Steps to Kolmogorov Complexity
$K_{\varepsilon}(\rho\mid R)$ $\leq C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\,\log\Gamma_{R}$ (32)
$\quad+c_{a}\,\log\!\Bigl(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}\Bigr)$
$\quad+c_{b}\,\log(1/\varepsilon)+O(1)$
@@
(Prefix Code Specification. A self-describing prefix code (such as Elias-delta) is used for $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ and $\varepsilon$. Therefore, there exist constants $c_{a},c_{b}=O(1)$ that depend only on the chosen universal machine and the prefix code, and are independent of $\rho,\Gamma_{R},d_{R}$, such that Eq. (32) holds.)